JEF meeting, May 2, 2025

From GlueXWiki
Jump to: navigation, search

Meeting time and location:

Agenda

  • ECAL maintenance and performance during the run
    • Issues with the HV [1] [2] [3]
    • Status of the calorimeter after the maintenance day [4]
    • Integrating ECAL into the trigger [5]
      • Run conditions and trigger rates [6]
    • HV module disconnected from the EPICS control
    • Bad flash ADC for reference PMTs
    • Low-level histograms [7]
    • Adjusting high voltages
      • iteration 5 (based on the run period 10, iteration 0) [8]
    • Recent ECAL occupancy, production at 300 nA with 4x10^-4 X0 Al radiator [9]. Suspicious modules after the accelerator maintenance day
    • Questions to address:
      • Rates in the detector
      • Gain stability
  • Calibrations
    • Initial HV settings using Compton scattering events [10]
    • Pi0 calibration page [11]
      • Pi0s are clearly seen for most ECAL modules (except in the innermost ring) after timing calibration.
  • Monitoring
  • Discussion of analysis topics and task division

Minutes

Attendees: Hakob, Igal, Olga, Liping, Phil, Sasha, Bill, Drew, Vlad, Yuyao, Zisis, S.T.

  • ECAL detector has been commissioned and incorporated into the trigger
    • 5 iterations for hardware HV settings
    • Gains ~1 with 2% width
    • Issues with HV: 2 modules dropped HV, failed twice in a day
      • External power cycling of crates now available
    • 2 connector problems identified, both fixed but one does not have a final fix (needs connector replacement)
    • Need to study performance at high rate and monitor gain stability over time
      • Look at LMS data for stability
    • Temperature dependence: saw shift in π0 mass position correlated with temperature
    • 4 inner rings excluded from trigger
      • Need to exclude more rings to go to higher rates
  • Igal working to add ECAL to π0 skim
  • Discussion of JEF physics priorities
    1. π0γγ Standard Model measurement then look for Beyond Standard Mode physics
    2. π+π-π0 Dalitz analysis then mirror asymmetry + π-π+ resonance search (BSM)
    3. e+e-γ dark photon search
    4. 2γ normalization reaction + Primakoff
    • Would like to have preliminary result in time to submit proposal to next PAC for more time
      • Need to understand backgrounds
      • Will not have final calibrations in place
        • Calibrations are ongoing procedure as data is accumulated -- use student help for initial calibrations?
        • Need good CDC calibrations for recoil proton, can consider not requiring reconstruction of proton: missing mass technique
      • For Primakoff, special Compton runs? How much data do we need? or triplet production?
  • Small scale prototype to be installed in PS arm
    • Can move to hall
    • Anticipate data taking in June
    • Student identified